Male Group (n=16, mean=11.00, SD=4.00) vs Female Group (16 females; mean=15.00, Variance = 9.00). Calculate the standard error of mean
for the male and female group. Show your work with your answer.

44. Calculate standard error of distance between male and female group on the basketball free-throws. Show your work with your answer.

-. Calculate the t-score for the basketball free-throws between male and female groups.

-. Report the t test result in a statistical term, including the t-score and the significant level.

-. Using the t-table below, the t value with the degree of freedom for this analysis (i.e., your answer on Question #1) should be equal or great
than ______________________________________ at the p value of 0.05 (95%) with two-tails.
-. In conclusion, is there any gender difference on free throws? Yes or no, and why? Do you accept or reject the null hypothesis?
Answer #1

Null hypothesis : means are equal

Alternative hypothesis : means are different

This is a two tailed unpaired t test with unequal variances.

his is Unpaired t-tet example with He re unequal variances 2 2 22 2 Hhe values Iutteg q2 41 have not here, n,&ng an df iven bputting values of both sample sizes

We get degree of freedom = df = 1.38×(16-1) = 20.7 = 21

Standard error of male group = 4/4 = 1

Standard error of female group = 9/4 = 2.25

Standard error of difference of two means = 2.462

t score = (11-15)/ 2.462 = -1.6247

Critical values from t table as per df (=21) and alpha (=0.05) is 2.08 which is greater than t- statistic.

Hence , we fail to reject the null hypothesis.

i.e we accept the null hypothesis.

No,There is no significant evidence to support the claim that the two means are different.